7. HHH Inc. reported $32,000 of sales and $8,700 of operating costs (including depreciation). The company had $16,000 of investor-supplied operating assets (or capital), the weighted average cost of that capital (the WACC) was 12.5%, and the federal-plus-state income tax rate was 40%. What was HHH's Economic Value Added (EVA), i.e., how much value did management add to stockholders' wealth during the year? a. $14,256 b. $13,100 c. $12,005 d. $11,980

  1. Arshadi Corp.'s sales last year were $52,000, and its total assets were $22,000. What was its total assets turnover ratio (TATO)?

           

  1.          2.364
  2.          2.198
  3.          1.695
  4.          1.263
  1. Rappaport Corp.'s sales last year were $350,000, and its net income after taxes was $26,000. What was its profit margin on sales?

           

  1.          9.558%
  2.          8.065%
  3.          7.429%
  4.          6.664%
Answer #1

1.
=(Sales-Operating costs)*(1-tax rate)-Invested Capital*WACC
=(32000-8700)*(1-40%)-16000*12.5%
=11980

2.
=Sales/Total Assets
=52000/22000
=2.363636364

3.
=Net Income/Sales
=26000/350000
=7.429%